OpenStudy (anonymous):

Please help dont understand... Solve the equation. √ 3x-5=4 7 3 1/3 49

OpenStudy (anonymous):

The Seven is one of the options. Sorry.

OpenStudy (anonymous):

Is the root aver the 3x only?

OpenStudy (anonymous):

only over the 3x*

OpenStudy (anonymous):

until the five.

OpenStudy (anonymous):

\[\sqrt{3x-5}=4\]

OpenStudy (anonymous):

If you could explain too, I would really appreciate that.

OpenStudy (anonymous):

First square both sides to get rid of the square root: 3x-5 = 4^2 Then subtract 5 from both sides to isolate 3x: 3x - 5 +5 = 16 + 5 Divide both sides by 3 to isolate x: 3x/3 = 21/3 So simplifying x = 7

OpenStudy (anonymous):

Cool. Thnx @znimon

OpenStudy (anonymous):

Any questions? just let me know

OpenStudy (anonymous):

Does this work if theres a number outside/aka in front of the √ symbol

OpenStudy (anonymous):

Do you mean \[2 \times \sqrt{somthing}\] or \[\sqrt[5]{somthing}\]

OpenStudy (anonymous):

\[4\sqrt{x-5-8}=12\]

OpenStudy (anonymous):

When solving an equations you are following reverse order or operations to essentially undo everything that has been done to x. So, in the case you just mentioned you would divide both sides by 4 then square both sides

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

These are the options for it... 20 30 25 25/4

OpenStudy (anonymous):

we would be dividing the 4 and the 12 right?

OpenStudy (anonymous):

But it turns out that if you were to square both sides in this case it wouldn't matter. I don't want to confuse you and it is a good habit to get into to always use reverse order of operations. You would get \[(4\sqrt{x-5-8})^{2} = 12^2\]

OpenStudy (anonymous):

we dont divide, just square?

OpenStudy (anonymous):

simplify inside the square root first

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

x-5-8 right?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

haha - 13 sorry

OpenStudy (anonymous):

\[4 * \sqrt{x-13} = 12\]

OpenStudy (anonymous):

Then divide both sides by 4: \[\sqrt{x-13} = 3\] Square both sides: \[x - 13 = 9\] Add 13 to both sides: x = 22

OpenStudy (anonymous):

22 isn't an option though...

OpenStudy (anonymous):

x = 22 is the correct answer

OpenStudy (anonymous):

Do we round?

OpenStudy (anonymous):

Well do the instructions say to round?

OpenStudy (anonymous):

20 is the closest option to 22 so...maybe 20?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[4 * \sqrt{x-5-8} = 12\] ?

OpenStudy (anonymous):

why the question mark? Is there something your unsure of?

OpenStudy (anonymous):

I am checking is if this is the right problem because the only solution for x is 22

OpenStudy (anonymous):

yes this is right.

OpenStudy (anonymous):

I'm not sure perhaps the possible answers are faulty the solution in x = 22

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

ok, well we tried our best, can you help me with this one too please, thanks again. \[\sqrt{6m-5}=m\]

OpenStudy (anonymous):

@znimon

OpenStudy (anonymous):

Try this one on your own and ask me questions as you go. Our previous solution to the previous problem is absolutely correct.

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

ok i'll try.

OpenStudy (anonymous):

do i divide by six?

OpenStudy (anonymous):

you need to square both sides first

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

How to write exponent?@znimon

OpenStudy (anonymous):

@Callisto